find the chi-squared value x2L corresponding to a sample size of 7 and a confidence level of 99%
Answer #1

Solution :

Given that,

n = 7

df = 7 - 1 = 6

2L = 2/2,df = 18.548

2R = 21 - /2,df = 0.676